OpenStudy (anonymous):

So say that the sum of the interior angle is 4320 degrees. How would you find the degree of each angle and the number of sides of the polygon?

sam (.sam.):

Use the formula: \[\frac{(n-2)*180}{n}=4320\] Find n

OpenStudy (anonymous):

So you need to divide it by n in order to find the degrees of each angle?

sam (.sam.):

Wait, my equation is finding ONE angle,so, (n-2)*180=4320 Find n

sam (.sam.):

n should be =26

OpenStudy (anonymous):

n is finding the number of sides

sam (.sam.):

once you have n=26. \[\frac{(26-2)180}{26}=166.15\]

sam (.sam.):

yes,n= number of sides

OpenStudy (anonymous):

what about finding the degree(s) of each interior angle?

sam (.sam.):

166.15, Found it on top

OpenStudy (anonymous):

Oh, I see!

Directrix (directrix):

If the polygon is not regular, you cannot determine the measure of each interior angle.